RC Meeting and Run Plan, Monday, February 24th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

Friday 24th Feb. 2020.

General updates

  • Relatively smooth weekend with the following exceptions:
    • Lost beam for ~7 1/2 h during the day on Saturday due initially to an IOC problem, then a rash of NL problems & chopper issue.
    • AN hour or so before this down, the target pol (L) was 55%, and it was changed to transverse for a planned 5 h run.
    • When it became apparent the down was going to be a longer one, Xiaochao determined we had enough transverse in the bag and the target was changed back to longitudinal, and the spectrometer kinematics were changed back to the HMS/SHMS -3.5/-3.4 GeV momentum settings.
    • Five h after the previous NMR, it was measured again and had dropped to 43%!!! https://logbooks.jlab.org/entry/3789864
    • When beam finally came back, we resumed running at the -3.5/-3.4 kinematics, Longitudinal pol. The next NMR was 50% (recovering well) and eventually the tgt pol spun back up to ~55%.
    • Then again this morning, we went from 55% just after midnight, to 49.6% at 5:47, and 50.8% at 10:59. SO we lost 5% but it's recovering.
    • The target experts are trying to figure out why this is happening. We should probably not run transverse again until we understand what's going on.
  • Friday night we had another storm of mm-scale vertical excursions on our 3H07 BPMs. We had pointed out to the PD and the crew chief on Friday morning that this was a problem for us and they were watching for it. The crew chief that evening saw it and managed to correlate it with the Hall A FFB RF going off & on. He made some changes in the accelerator (see https://logbooks.jlab.org/entry/3789212 and https://logbooks.jlab.org/entry/3789222). We have had some occasional 0.3-0.4 mm excursions since then, but no big ones.
  • Tomorrow the beam goes off 8:30 -12:30 AM (could be a little longer). We will access the hall opportunistically:
    • The techs (4-5 keys) have 3 pumps lined up to pump the xfer line IV on the HMS & SHMS, as well as the Moller magnet IV.
    • Cryo (1 key) will come in with a sonic sniffer to look for the source of the 300 l/d warm He leak they have had since the ESR crash last Thursday. They will check other halls too, and the ESR itself today, but they're worried it might be the Moller warm return line again, even though it was repaired.
    • JP said the target experts may want to work on the target (laser access mode) at some point to investigate the polarization drops.
  • Moller measurement Thursday:
    • Assuming there are no leaks discovered tomorrow in the Moller warm return, we'll pre-cool the Moller lines overnight Tuesday
    • Cool-down the Moller Wednesday morning, or if cryo is part of the safety standown, Wed afternoon. We have coordinated with them on this already.
    • Do the Moller measurement Thursday day
  • Three hour-long safety stand-downown for accelerator Wednesday morning 9-12 AM (recent rumors suggest it may include engineering too)
    • No access to the hall unless we were to go to restricted access much earlier than 9 and deal with the baggage to return to beam permit afterwards. We do not plan to do that. We will stay in power permit along with the accelerator.
    • No details of what the stand-down will address, although there have been almost a dozen major safety incidents by my count since the beginning of the year. More than 1 per week on average. Two just today, and the day isn't over yet.

Pending problems

  • 3 hour stand-down Wed AM.
  • Locks on vertical beam position have not been working (see https://logbooks.jlab.org/entry/3788540) but seem better since last Friday night's adjustments (see above).
  • Small changes in beam position have been deliberately made as well while we are at 30 uA on 3He (see https://logbooks.jlab.org/entry/3788864)
  • Raster energy alarm- this should have been fixed but I have forgotten to followup on it.
  • Harp travel problem can't be fixed without taking the harps down (out). Limit switch is already at end of clearance. Way bigger job than we thought (so punt).
    • In the meantime, since we take our harp scans with tune beam, we have been asking to deliberately move the beam 1-2 mm (1 mm is just barely enough) for the harp scan, then to move it back. We only care about the widths so this seems like a reasonable work-around for now.

Not sure if it is a problem

Fixed problems

  • Brad added the Orbit lock indicator to our Hall C wall screen, making it easier to check if it's on or not
  • Friday night adjustments by the crew chief (see above) may have ended the larger >1mm vertical excursions we have been having. But the jury's still out.
  • Lots of noise on our energy since Hall A stole our 2A lock, and have been trying to get FFB turned on
    • Since these are 2-3 MeV, ie dp/p ~ 2e-4 it's within specs and shouldn't impact our physics. So I haven't been complaining about it.
  • Trip rate is still high, averaging ~15 trips/h most of the time

Opportunistic Problems to get done at some point

  • Note there may need to be a recesiation of the photocathode in 10-14 d which could take 10-12 h, according to what Joe Grames said at Wednesday's 1:30 meeting. See https://logbooks.jlab.org/entry/3788008
    • The injector group may just move the spot- it remains to be seen.
    • This could be checked when we go down for 4 h Tuesday to fix a PSS problem with the lasers in the injector
    • Our QE is still reasonable, ~0.29 this AM (see https://logbooks.jlab.org/entry/3790779), was 0.32 a few days ago though

Opportunistic Access Jobs

  • See above

Recent polarization measurements

Current Run Plan

  • Taking production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization until further notice. NMR every 5 hrs.
  • Tuesday when we come out of the access, we'll take a 1h run with target out in order to measure the background from just the Beryllium windows.
  • Next Moller measurement Thursday.
  • If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change then beam tuning (except work on transport magnets), then do beam recovery procedure (do YAG viewer part only if you are suspicious of the harp scan results)
  • Check the position locks are on.
  • Check CODA is running normally (had one strange hangup lately)
  • CHeck online replays shortly after starting runs, & full replay results too.
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.

!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!


  • The transition to d2n will probably happen sometime the week of March 9-13.
  • For the time being we will run exclusively with longitudinal polarization. Do NMR before and after transverse, and every 5 hours during longitudinal.
  • Production at 30 degrees, HMS/SHMS=-3.5/-3.4 momentum.

When there is a down:

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If SHMS dipole trips again, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]
  • Do not change SHMS momentum until we fixed the current lead valve

Planned Accelerator Configuration Changes

  • Note there will likely be a recesiation of the photocathode in 10-14 d which could take 10-12 h, according to what Joe Grames reported at the 2/19 1:30 meeting.
    • The injector group may just move the spot- it remains to be seen.
    • This could be checked when we go down for 2-3 h in the next few days to fix a PSS problem with the lasers in the injector
  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links